A stone is thrown downwards from the top of the castle wall with an initial speed 20.0m/s. The wall is 30m high
a) what is the stone's speed just as it arrives at ground level?
b) How much time between throwing and landing?
c) sketch position, speed and accelerarion graphs to show what happened
1
Expert's answer
2019-10-28T10:55:28-0400

Fail time is the positive root of the equation: "t^2 + 20t - 6 = 0", it is 1.16 s.

Accordingly, the velocity at the time of the fall v = 20 + 11.6 = 31.6 m/s.